40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 17154     Entry time: Fri Sep 23 10:05:38 2022
Author: JC 
Type: Update 
Category: VAC 
Subject: N2 Interlocks triggered  

[Chub, Anchal, Tega, JC]

After replacing an empty tank this morning, I heard a hissing sound coming from the nozzle area. It turns out that this was from the copper tubing. The tubing we slightly broken and this was comfirmed with soapy water bubbles. This caused the N2 pressure to drop and the Vac interlocks to be triggered. Chub and I went ahead and replaced the fitting in connected this back to normal. Anchal and Tega have used the Vacuum StartUp procedures to restore the vacuum to normal operation.

Adding screenshot as the pressure is decreasing now.

Attachment 1: Screenshot_2022-09-23_10-19-39.png  41 kB  Uploaded Fri Sep 23 11:19:48 2022  | Hide | Hide all
Screenshot_2022-09-23_10-19-39.png
Attachment 2: Screenshot_2022-09-23_10-22-45.png  43 kB  Uploaded Fri Sep 23 11:22:57 2022  | Hide | Hide all
Screenshot_2022-09-23_10-22-45.png
ELOG V3.1.3-